OpenStudy (kiki99):

Solve this equation, step by step, using the methods mentioned below : 3x^2-12x-15=0 Factor:

OpenStudy (kiki99):

help i need step by step

OpenStudy (kiki99):

i started with adding +15 to both sides

OpenStudy (er.mohd.amir):

no use splitting middle term method or use quadratic formula

OpenStudy (kiki99):

can you show me im a little confused

OpenStudy (er.mohd.amir):

3x*x-(15-3)*x-15 =3x*x-15x+3x-15

OpenStudy (er.mohd.amir):

take common from first two terms and last two terms

OpenStudy (kiki99):

3?

OpenStudy (er.mohd.amir):

3x (x-5) +3 (x-5) =(x-5) (3x+3) =(3x+3) (x-5) =3(x+1) (x-5)